P rove that all real of the form a + b 2 , a; b ∈ Z form s a ring.


1
Expert's answer
2022-02-07T17:14:11-0500

We are to prove that the set "\displaystyle
S=\{a+b\sqrt{2}:a,b\in\Z\}" is a ring.

Let "p=a_1+b_1\sqrt{2},\ q=a_2+b_2\sqrt{2},\ r=a_3+b_3\sqrt{2}" belong to "S", then


1] "S"  is an abelian group under addition, meaning that:

"-(p+q)+r=p+(q+r)"

we have:

"(a_1+b_1\sqrt2+a_2+b_2\sqrt2)+a_3+b_3\sqrt2=a_1+b_1\sqrt2+(a_2+b_2\sqrt2+a_3+b_3\sqrt2)"


"--p+q=q+p"

we have:

"a_1+b_1\sqrt2+a_2+b_2\sqrt2=a_2+b_2\sqrt2+a_1+b_1\sqrt2"


"---" There is an element "0=0+0\sqrt{2}" in "S" if "a_1,\ b_1=0"  such that "p + 0 = p"


"----" For each p in S there exists −p in S such that p + (−p) = 0

we have:

"a_1+b_1\sqrt2+(-a_1-b_1\sqrt2)=0"


2] "S"  is a monoid under multiplication, meaning that:

"(p \u22c5 q) \u22c5 r = p \u22c5 (q \u22c5 r)"

we have:

"((a_1+b_1\sqrt2)(a_2+b_2\sqrt2))(a_3+b_3\sqrt2)=(a_1+b_1\sqrt2)((a_2+b_2\sqrt2)(a_3+b_3\sqrt2))"


There is an element "1=1+0\sqrt{2}" in "S\ \text{if }a=1, b=0"  such that p ⋅ 1 = p and 1 ⋅ p = p


3] Multiplication is distributive with respect to addition, meaning that:

p ⋅ (q + r) = (p ⋅ q) + (p ⋅ r)

we have:

"(a_1+b_1\sqrt2)(a_2+b_2\sqrt2+a_3+b_3\sqrt2)=(a_1+b_1\sqrt2)(a_2+b_2\sqrt2)+"

"+(a_1+b_1\sqrt2)(a_3+b_3\sqrt2)"


(q + r) ⋅ p = (q ⋅ p) + (r ⋅ p)

we have:

"(a_2+b_2\sqrt2+a_3+b_3\sqrt2)(a_1+b_1\sqrt2)=(a_1+b_1\sqrt2)(a_2+b_2\sqrt2)+"

"+(a_1+b_1\sqrt2)(a_3+b_3\sqrt2)"